statistics grades: In a statistics class of 35 students, there were 17 men and 18 women. Six of the men and seven of the women received an A in the course. A student is chosen at random from the class. (a) Find the probability that the student is a woman. (b) Find the probability that the student received an A. (c) Find the probability that the student is a woman or received an A. (d) Find the probability that the student did not receive an A.
Answer #1
Men Women Total
Receive A 6 7 13
Not receive A 11 11 22
Total 17 18 35

a)

P(Women) = 18/35

= 0.5143

b)

P(A) = 13/35

= 0.3714

c)

P(Women OR A) = P(Women ) + P(A) - P(Women and A )

= 18/35 + 13/35 - 7/35

= 24/35

= 0.6857

d)

P(not A) =22/35

= 0.6286